# Seeing Beyond the Image: ECG and Anatomical Knowledge-Guided Myocardial Scar Segmentation from Late Gadolinium-Enhanced Images

### v1

@@ Abstract
Accurate segmentation of myocardial scar from late gadolinium enhanced (LGE) cardiac MRI is essential for evaluating tissue viability, yet remains challenging due to variable contrast and imaging artifacts. Electrocardiogram (ECG) signals provide complementary physiological information, as conduction abnormalities can help localize or suggest scarred myocardial regions. In this work, we propose a novel multimodal framework that integrates ECG-derived electrophysiological information with anatomical priors from the AHA-17 atlas for physiologically consistent LGE-based scar segmentation. As ECGs and LGE-MRIs are not acquired simultaneously, we introduce a Temporal Aware Feature Fusion (TAFF) mechanism that dynamically weights and fuses features based on their acquisition time difference. Our method was evaluated on a clinical dataset and achieved substantial gains over the state-of-the-art image-only baseline (nnU-Net), increasing the average Dice score for scars from 0.6149 to 0.8463 and achieving high performance in both precision (0.9115) and sensitivity (0.9043). These results show that integrating physiological and anatomical knowledge allows the model to “see beyond the image”, setting a new direction for robust and physiologically grounded cardiac scar segmentation.
@@ 4 Results and Analysis
Our multi-modal approach significantly outperforms the image-only baseline. Using only LGE-MRI, the baseline achieves a Dice of 0.6149, reflecting the challenges of scar segmentation under contrast variability and artifacts. Adding ECG features, anatomical priors, and time-aware fusion yields a markedly higher Dice score of 0.8463, demonstrating a substantial improvement over 37%. This demonstrates how physiological and anatomical information improves scar localization and boundary accuracy beyond image cues alone, as shown in Figure 2. Ablation results are also reported in Table 1.
@@
[0,3)[3,7)[7,14)[14,21)[21,30)[30,60)[60,90]0.70.70.80.80.90.9Time interval (days) binsMean Dicew/o time conditioningw/ time conditioning
Fig. 3: Mean Dice across time-interval bins for models w/ and w/o time conditioning.
Result visualization and interpretability. Figure 2 compares segmentation quality (b,c) and region-wise scar volume across 17 AHA myocardial regions for the ground truth (d), the image-only baseline prediction (e), and our multimodal prediction (f) for a test subject. The ground truth plot shows scar burden in specific mid-ventricular and apical regions. The baseline model captures the broad pattern but consistently underestimates scar, while the multimodal model more closely matches both the magnitude and spatial distribution across the 17 myocardial regions. This reflects the added value of ECG and temporal cues in capturing subtle regional variations in scar distribution, resulting in uniform performance. The temporal attention maps (h) highlight that the model focuses more on the clinically indicative regions/biomarkers such as Q waves, QRS and/or abnormal ST–T segments [Das2008-oq, Kashou2025-dy], on leads II, aVR, V2, and V4 and the cross-lead attention map (g) highlights the distinctive pattern in Lead III and shows a positive correlation among V5, V6, and Lead II.
Limitations and future work. The ECG-driven gains are promising but remain incremental, likely constrained by the quality and limited information from digitized paper ECGs. Future gains may be achieved by directly using full 10-second digital ECGs or developing more robust digitization algorithms as well as designing better temporal-aware feature fusion. As shown in Fig. 3, the model with time conditioning achieved more significant improvements on cases with a 1-week [7,14) and 3-week to 1-month time window compared to the rest. Our current segmentation remains 2D, so extending it to 3D to better capture heart structure and function can be another promising direction.

### v2

@@ Abstract
Accurate segmentation of myocardial scar from late gadolinium enhanced (LGE) cardiac MRI is essential for evaluating tissue viability, yet remains challenging due to variable contrast and imaging artifacts. Electrocardiogram (ECG) signals provide complementary physiological information, as conduction abnormalities can help localize or suggest scarred myocardial regions. In this work, we propose a novel multimodal framework that integrates ECG-derived electrophysiological information with anatomical priors from the AHA-17 atlas for physiologically consistent LGE-based scar segmentation. As ECGs and LGE-MRIs are not acquired simultaneously, we introduce a Temporal Aware Feature Fusion (TAFF) mechanism that dynamically weights and fuses features based on their acquisition time difference. Our method was evaluated on a clinical dataset and achieved substantial gains over the state-of-the-art image-only baseline (nnU-Net), increasing the average Dice score for scars from 0.6149 to 0.8463 and achieving high performance in both precision (0.9115) and sensitivity (0.9043). These results show that integrating physiological and anatomical knowledge allows the model to “see beyond the image”, setting a new direction for robust and physiologically grounded cardiac scar segmentation. Code is available at this URL.
@@ 4 Results and Analysis
Our multi-modal approach significantly outperforms the image-only baseline. Using only LGE-MRI, the baseline achieves a Dice of 0.6149, reflecting the challenges of scar segmentation under contrast variability and artifacts. Adding ECG features, anatomical priors, and time-aware fusion yields a markedly higher Dice score of 0.8463, demonstrating a substantial improvement over 37%. This demonstrates how physiological and anatomical information improves scar localization and boundary accuracy beyond image cues alone, as shown in Figure 2. Ablation results are in Table 1.
@@
Fig. 3: Distribution of Dice scores across time-interval bins, comparing models with and without time conditioning.
Result visualization and interpretability. Figure 2 compares segmentation quality (b,c) and region-wise scar volume across 17 AHA myocardial regions for the ground truth (d), the image-only baseline prediction (e), and our multimodal prediction (f) for a test subject. The ground truth plot shows scar burden in specific mid-ventricular and apical regions. The baseline model captures the broad pattern but consistently underestimates scar, while the multimodal model more closely matches both the magnitude and spatial distribution across the 17 myocardial regions. This reflects the added value of ECG and temporal cues in capturing subtle regional variations in scar distribution, resulting in uniform performance. The temporal attention maps (h) highlight that the model focuses more on the clinically indicative regions/biomarkers such as Q waves, QRS and/or abnormal ST–T segments [Das2008-oq, Kashou2025-dy], on leads II, aVR, V2, and V4. The cross-lead attention map (g) highlights the distinctive pattern in Lead III and shows a positive correlation among V5, V6, and Lead II.
Limitations and future work. The ECG-driven gains are encouraging but remain modest, likely due to variability in digitalized ECG signal quality and the imbalanced sample distribution across bins. Future gains may be achieved by developing more robust digitization algorithms as well as designing better temporal-aware feature fusion with larger datasets and extending 2D to 3D. As shown in Fig. 3, the model with time conditioning achieved more significant improvements on cases with a 1-week [7,14) and 3-week to 1-month time window compared to the rest. The time-conditioned model shows higher median performance across most of the time bins, with more pronounced gains at longer time intervals indicating improved robustness to large temporal gaps between modalities.